# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

### v1

@@ Abstract
Despite significant advancements in text generation and reasoning, Large Language Models (LLMs) still face challenges in accurately performing complex arithmetic operations. To achieve accurate calculations, language model systems often enable LLMs to generate code for arithmetic operations. However, this approach compromises speed and security and, if finetuning is involved, risks the language model losing prior capabilities. We propose a framework that enables exact arithmetic in a single autoregressive step, providing faster, more secure, and more interpretable LLM systems with arithmetic capabilities. We use the hidden states of an LLM to control a symbolic architecture which performs arithmetic. Our implementation using Llama 3 8B Instruct with OccamNet as a symbolic model (OccamLlama) achieves 100% accuracy on single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), outperforming GPT 4o and on par with GPT 4o using a code interpreter. OccamLlama also outperforms both Llama 3 8B Instruct and GPT 3.5 Turbo on multistep reasoning problems involving challenging arithmetic, thus enabling small LLMs to match the arithmetic performance of even much larger models. We will make our code public shortly.
@@ 1 Introduction
Since the release of GPT 3, Large Language Models (LLMs) have dramatically improved in their text generation and reasoning capabilities. This has enabled success in downstream applications including machine translation [1, 2], sentiment analysis [3, 4, 5], and interactive dialogue generation [6], with language models even surpassing human experts on some academic benchmarks that require reading comprehension, reasoning and coding [7]. However even industry-leading LLMs such as GPT 4 cannot reach 100% accuracy on simple arithmetic [8], limiting their ability to perform basic mathematical tasks. This hinders potential applications of LLMs ranging from chat-bot physics tutors to LLM-powered automated research which would accelerate scientific discovery and technological innovation. The poor arithmetic performance of LLMs is particularly acute for small LLM agents, limiting their usage in smartphone or in multi-agent applications.
To enable accurate calculations, language model systems often resort to running code written by an LLM. However, this comes at the cost of speed; the model must perform multiple autoregressive steps to generate code that performs the appropriate arithmetic operations. This increased decoding time may negatively impact applications such as multi-agent workflows [9, 10] where speed is essential. At the same time, code-based LLM arithmetic mechanisms may increase system vulnerability by providing a mechanism for arbitrary LLM-generated code execution.
We propose an alternative, a framework which enables exact and interpretable LLM arithmetic in a single autoregressive step, providing faster and more secure arithmetic capabilities in LLM systems. Our framework uses the hidden states of an LLM to control a symbolic architecture that performs arithmetic. Although our method can in principle work with any symbolic architecture, in this paper we use an interpretable neurosymbolic architecture known as OccamNet [11, 12] because of its interpretability and scalability. Because of this, we term our method OccamLLM, or OccamLlama when using a Llama model as the LLM.
@@
    We benchmark OccamLlama on arithmetic tasks, demonstrating that OccamLlama achieves 100% accuracy on arbitrary single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), more than double the accuracy of GPT 4o. OccamLlama performs slightly better than GPT 4o with Code Interpreter while using on average more than 50x fewer tokens to generate.
@@
    We benchmark on mathematical reasoning problems, showing that OccamLlama can sustain long generations. OccamLlama outperforms both Llama 3 8B Instruct and GPT 3.5 Turbo on reasoning problems with challenging arithmetic.
@@ 2 Related Work
Arithmetic Performance in LLMs.
Prior research has trained models on synthetic data, finding that such models can achieve near-perfect accuracy on addition [14, 15], subtraction [15], multiplication [14, 15], division [15], and raising to powers [15]. Furthermore, these prior models have been tested only on arithmetic datasets, so their generality has not been assessed.
Other work focuses on finetuning LLMs which are already trained on large amounts of general-purpose data on math datasets. Both full-parameter [16, 17] and parameter-efficient (PEFT) [18] finetuning strategies have been applied in this regard. However, finetuning on a single dataset carries the risk of catastrophic forgetting of an LLM’s previously acquired linguistic skills [19]. While PEFT techniques have been shown to partially mitigate this effect, this area is still one of active research [20, 21].
LLMs with Tool Use.
Another thrust of prior research has focused on LLM tool use, which we believe is most directly related to our methods. Calc-X [22] introduces a technique to offload arithmetic computations to an external tool like a calculator. The authors curated a large dataset of arithmetic problems and trained a language model that learns to interact with a calculator through the use of tags to signify the calling of the external tool. Several other works [23, 24, 25] follow a similar idea, using crowd workers to annotate tool calls and using this data to train language models to interact with external tools such as a web searching tool, a calculator, or a translation system. These approaches can be prohibitively expensive in annotation costs; Toolformer [26] overcomes this cost by using in-context learning and a language model to generate datasets containing the necessary ‘API’ tool calls via a self-supervised loss. Further, the above methods all require finetuning of the LLM, placing the LLM at risk of losing generality and its original language modelling abilities through catastrophic forgetting. In contrast, our approach does not involve training the language model. Our ‘external tool’ is a symbolic neural network which can be trained to correctly use the hidden states of the language model to perform the required arithmetic computations. The language model is kept frozen throughout this process. Unlike other tool-calling approaches, where the cost of data annotation to train for tool-calling interaction can be prohibitively expensive, in our method, each task only requires manually annotating tens of prompts, a high annotation efficiency. Other prior methods leverage prompt engineering to improve arithmetic performance of LLMs; this is done either through chain-of-thought [27], or to encourage LLMs to use a code interpreter [28, 29, 30]. Contrary to these methods, our approach does not use code kernels; this provides several advantages: 1) it enables tool use without expending compute on autoregressive steps for token generation, and 2) it avoids running potentially incorrect or malicious code generated by language models.
@@ 3.1 OccamLLM: Combining a Language Model with a Symbolic Model
In short, the OccamLLM system combines a language model with a symbolic model, namely OccamNet, that can perform arithmetic operations like addition and subtraction. For each token, the corresponding internal hidden states of the language model are fed into a decoder module which initializes the symbolic model so that it executes the operation required by the task described in the input text. A string parser feeds the necessary numbers from the text into OccamNet, which evaluates the desired expression. Finally, a decoder determines whether to use the language model output or the OccamNet output for generating the next token.
In the example shown in Figure 1, a decoder determines how to initialize OccamNet from the language model hidden states, choosing to have OccamNet perform addition. The text parser then feeds the numbers $6$ and $7$ into the symbolic model, which adds the numbers, returning $13$ Finally, a decoder decides to use the OccamNet output instead of the language model output, so the system outputs 13. The new sentence, including the 13, is tokenized and fed back to the LLM to continue normal autoregressive generation. The language model might later generate “Since she ate two apples, she now has,” at which point the switch will again trigger OccamNet, this time implementing $13-2$ and returning $11$
In the subsections below we describe the OccamLLM system which we find performs substantially better than other LLMs. For a much larger analysis of alternate architectures and losses, see Appendix B.
@@ 3.1.1 OccamNet
OccamNet is a symbolic architecture that provides an interpretable way of parametrizing probability distributions over a space of functions [11]. We leave a more thorough explanation of OccamNet to [11], describing only the relevant components here. For a more detailed explanation, see Appendix C.
@@
OccamNet has the structure of an $n$-input, $l$-internal-activation-layer multilayer perceptron with the biases removed and the activations in each layer replaced by the primitives $\mathcal{P},$ as shown in Figure 2a. Activation functions may have multiple inputs. We rename the linear layers softmax layers, denote the weights of the $i$th softmax layer as $\mathbf{W}_{i}$, and denote the combined weights of OccamNet as $\mathbf{W}.$
@@
In principle, OccamLLM can work with any symbolic model, i.e., any model that can parameterize a set of symbolic functions or a distribution over such functions. We choose OccamNet as opposed to, for example, a transformer [31] or recurrent neural network [32], for two reasons: 1) OccamNet is interpretable, which we hypothesize makes controlling OccamNet easy for a decoder to learn, and 2) OccamNet is parallelizable over multiple samples, allowing for scalable training.
@@ 3.1.2 OccamLLM Decoder
The OccamLLM decoder takes the hidden states of a language model and outputs an initialization for OccamNet. This gives the LLM control over which function to apply on the inputs. The decoder acts on each input token separately, producing a different OccamNet initialization for each. Therefore, the arithmetic operations predicted may change along an input sequence, allowing OccamNet’s use for different computations in a single example. This is crucial in multi-step reasoning scenarios where OccamNet is employed several times for different purposes.
Many decoder architectures are possible. We choose to parameterize the weights of each softmax layer of OccamNet independently, as $(\mathbf{W}_{1},\ldots,\mathbf{W}_{l})=(\text{Decoder}_{1}(\mathbf{h}),\ldots,\text{Decoder}_{l}(\mathbf{h})),$ where $\mathbf{h}$ are the hidden states of the language model. We choose
$\text{Decoder}_{i}(\mathbf{h})=\text{MLP}_{i}\left(\sum_{j=1}^{L}w_{i,j}\mathbf{h}_{j}\right)+\mathbf{W}_{i}^{(0)}$ (1)
where $\mathbf{h}_{j}$ are the hidden states of the $j$th layer of the language model, $w_{i,j}$ are trainable weights, $\text{MLP}_{i}$ are two-layer multilayer perceptrons (MLPs), and $\mathbf{W}_{i}^{(0)}$ are untrained weights which initialize all functions to have approximately equal probabilities according to the initialization described in [11] and explained in Appendix C.4.
@@ 3.2 Data Generation
We create synthetic datasets to train the OccamLLM decoder, which contain instruction prompts for diverse arithmetic tasks. To generate datasets of arbitrary size, we create prompts with placeholders for numbers to sample from. Each prompt includes a string with number placeholders, the sampling value range for each number, and a function that computes the answer to the query given the sampled input numbers. The prompts fall into two main categories: purely arithmetic tasks and reasoning problems.
@@ 3.2.1 OccamNet Decoder Training Data
For training the decoder that controls the weights of OccamNet, we created two types of examples, single queries and concatenated queries. For single queries, we select a single prompt from the problems generated as discussed in Section 3.2. We use the Llama 3 8B Instruct chat template and fill in the query as the user input and the result as the assistant response, prepending “Answer = ” to the later in randomly selected samples (see Appendix A.1 for further details). For the concatenated queries of examples, we select a random number of prompts and concatenate the query-response pairs without using the Llama 3 8B Instruct chat template. The OccamNet decoder is trained to predict only the results of the last query in the sequence. This strategy helps OccamLLM to learn which operation to perform without becoming confused by the text that appears before, which is useful for continuous generation. To create the training dataset, each example is sampled by first randomly selecting the kind of example, then randomly selecting the type(s) of prompt(s) used, and finally randomly sampling the input values from the range corresponding to each selected prompt.
@@ 3.2.2 OccamLLM Switch Training Data
To train the switch, we generate examples of possible LLM outputs for given input expressions, and label the outputs with sequences of 0s or 1s corresponding to whether the language model or the OccamNet output should be used for the next token. Some input-output pairs correspond to the prompts described in Section 3.2. For this, the LLM output is set to “The answer is” or “Answer = ” and the label sequence only contains a 1 at the end to indicate the system should use OccamNet to compute the answer. We also manually created and labeled several other examples for diverse scenarios, to explicitly teach the system in which cases it should or should not use OccamNet. To create the dataset, we concatenate a random number of these user input - assistant output pairs in a conversational fashion, using the Llama 3 8B Instruct chat template.
@@ 3.3 OccamLLM Training
We train the OccamLLM decoder and the switch separately, as they do not share weights. In all cases, the weights of the LLM are kept frozen. In the first step, we train the system to predict the answer to examples generated by the method explained in Section 3.2.1. The OccamNet decoder needs to process the set of hidden states corresponding to the last token of the response and set the weights of OccamNet appropriately, so that the correct arithmetic expression is sampled. In this step, we use a rescaled REINFORCE [34] loss:
@@
where $p_{W}[f]\equiv\text{ON}(f;\text{Decoder}_{W}(\mathbf{h}(x)))$ is the probability distribution represented by the decoder-initialized OccamNet. We find setting $R(f(x),y)=1$ if $f(x)=y$, and 0 otherwise, most effective.
Minimizing this loss steers the decoder towards assigning higher probabilities to the functions that maximize the reward $R(f(x),y)$, which measures the similarity between the correct answer $y$ and the prediction of OccamNet $f(x)$. We discuss the OccamNet loss in more detail in Appendix B.
@@ 4 Experiments
This single layer OccamNet can be invoked by the LLM several times during generation to perform any arithmetic operation accurately. To use the trained OccamLlama for inference, we sample the highest probability function from OccamNet as described in Appendix C.3.

### v2

@@ Abstract
Despite significant advancements in text generation and reasoning, Large Language Models (LLMs) still face challenges in accurately performing complex arithmetic operations. To achieve accurate calculations, language model systems often enable LLMs to generate code for arithmetic operations. However, this approach compromises speed and security and, if finetuning is involved, risks the language model losing prior capabilities. We propose a framework that enables exact arithmetic in a single autoregressive step, providing faster, more secure, and more interpretable LLM systems with arithmetic capabilities. We use the hidden states of an LLM to control a symbolic architecture which performs arithmetic. Our implementation using Llama 3 8B Instruct with OccamNet as a symbolic model (OccamLlama) achieves 100% accuracy on single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), outperforming GPT 4o and on par with GPT 4o using a code interpreter. OccamLlama also outperforms GPT 4o both with and without a code interpreter on mathematical problem solving benchmarks involving challenging arithmetic, thus enabling small LLMs to match the arithmetic performance of even much larger models. We will make our code public shortly.
@@ 1 Introduction
Since the release of GPT 3, Large Language Models (LLMs) have dramatically improved in their text generation and reasoning capabilities. This has enabled success in downstream applications including machine translation [1, 2], sentiment analysis [3, 4, 5], and interactive dialogue generation [6], with language models even surpassing human experts on some academic benchmarks that require reading comprehension, reasoning and coding [7]. However even industry-leading LLMs such as GPT 4 cannot reach 100% accuracy on simple arithmetic [8], limiting their ability to perform basic mathematical tasks. This hinders potential applications of LLMs ranging from chat-bot physics tutors to LLM-powered automated research that could accelerate scientific discovery and technological innovation. The poor arithmetic performance of LLMs is particularly acute for small LLM agents, limiting their usage in smartphone or in multi-agent applications.
To enable accurate calculations, language model systems often resort to running code written by a LLM. However, this comes at the cost of speed; the model must perform multiple autoregressive steps to generate code that performs the appropriate arithmetic operations. This increased decoding time may negatively impact applications such as multi-agent workflows [9, 10] where speed is essential. At the same time, code-based LLM arithmetic mechanisms may increase system vulnerability by providing a mechanism for arbitrary LLM-generated code execution.
We propose an alternative, a framework which enables exact and interpretable LLM arithmetic in a single autoregressive step, providing faster and more secure arithmetic capabilities in LLM systems. Our framework uses the hidden states of a LLM to control a symbolic architecture that performs arithmetic. Although our method can in principle work with any symbolic architecture, in this paper we use an interpretable neurosymbolic architecture known as OccamNet [11, 12] because of its interpretability and scalability. Therefore, we term our method OccamLLM, or OccamLlama when using a Llama model as the LLM.
@@
    We benchmark OccamLlama on arithmetic tasks, demonstrating that OccamLlama achieves 100% accuracy on arbitrary single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), more than double the accuracy of GPT 4o. OccamLlama performs slightly better than GPT 4o with Code Interpreter while answering in on average more than 50x fewer generation tokens.
@@
    We benchmark on mathematical problem solving tasks, showing that OccamLlama can sustain long generations. OccamLlama outperforms both GPT 4o and GPT 4o with code interpreter on benchmarks involving challenging arithmetic.
@@ 2 Related Work
Arithmetic Performance in LLMs.
Prior research has trained models on synthetic data, finding that such models can achieve near-perfect accuracy on addition [14, 15], subtraction [15], multiplication [14, 15], division [15], and raising to powers [15]. These prior models have been tested only on arithmetic datasets, so their generality has not been assessed. Other work focuses on finetuning LLMs which are already trained on large amounts of general-purpose data on math datasets. Both full-parameter [16, 17] and parameter-efficient (PEFT) [18] finetuning strategies have been applied. However, finetuning on a single dataset carries the risk of catastrophic forgetting of an LLM’s previously acquired linguistic skills [19]. While PEFT techniques have been shown to partially mitigate this effect, this area is still one of active research [20, 21].
LLMs with Tool Use.
Another thrust of prior research has focused on LLM tool use, which we believe is most directly related to our methods. Calc-X [22] introduces a technique to offload arithmetic computations to an external tool like a calculator. The authors curated a large dataset of arithmetic problems and trained a language model that learns to interact with a calculator through the use of tags to signify the calling of the external tool. Several other works [23, 24, 25] follow a similar idea, using crowd workers to annotate tool calls and using this data to train language models to interact with external tools such as a web searching tool, a calculator, or a translation system. These approaches can be prohibitively expensive in annotation costs; Toolformer [26] overcomes this cost by using in-context learning and a language model to generate datasets containing the necessary ‘API’ tool calls via a self-supervised loss. Further, the above methods all require finetuning of the LLM, placing the LLM at risk of losing generality and its original language modelling abilities through catastrophic forgetting. In contrast, our approach does not involve training the language model. Our ‘external tool’ is a symbolic model which can be trained to correctly use the hidden states of the language model to perform the required arithmetic computations. The language model is kept frozen throughout this process. Unlike other tool-calling approaches, where the cost of data annotation to train for tool-calling interaction can be prohibitively expensive, in our method, each task only requires manually annotating tens of prompts, a high annotation efficiency. Other prior methods leverage prompt engineering to improve arithmetic performance of LLMs; this is done either through chain-of-thought [27], or to encourage LLMs to use a code interpreter [28, 29, 30]. Contrary to these methods, our approach does not use code kernels; this provides several advantages: 1) it enables tool use without expending compute on autoregressive steps for token generation, and 2) it avoids running potentially incorrect or malicious code generated by language models.
@@ 3.1 OccamLLM: Combining a Language Model with a Symbolic Model
In short, the OccamLLM system combines a language model with a symbolic model, namely OccamNet, that can perform arithmetic operations like addition and subtraction. For each token, the corresponding internal hidden states of the language model are fed into a decoder module which initializes the symbolic model such that it executes the operation required by the task described in the input text. A string parser feeds the necessary numbers from the text into OccamNet, which evaluates the desired expression. Finally, a decoder determines whether to use the language model output or the OccamNet output for generating the next token.
In the example shown in Figure 1, a decoder determines how to initialize OccamNet from the language model hidden states, choosing to have OccamNet perform addition. The text parser then feeds the numbers $6$ and $7$ into OccamNet, which adds the numbers, returning $13$ Finally, a decoder decides to use the OccamNet output instead of the language model output, so the system outputs 13. The new sentence, including the 13, is tokenized and fed back to the LLM to continue autoregressive generation. The language model might later generate “Since she ate two apples, she now has,” at which point the switch will again trigger OccamNet, this time implementing $13-2$ and returning $11$
In the subsections below, we describe the OccamLLM system which from our experiments we find to be most performant, even oupterforming GPT 4o in several benchmarks. For an analysis of alternate architectures and losses, see Appendix C.
@@ 3.1.1 OccamNet
OccamNet is a symbolic architecture that provides an interpretable way of parametrizing probability distributions over a space of functions [11]. We leave a more thorough explanation of OccamNet to [11] and Appendix D, describing only the relevant components here.
@@
OccamNet has the structure of an $n$-input, $l$-internal-activation-layer multilayer perceptron with the biases removed and the activations in each layer replaced by the primitives $\mathcal{P},$ as shown in Figure 2a. Activation functions may have multiple inputs. We rename the linear layers softmax layers, denote the weights of the $i$th softmax layer as $\mathbf{W}^{(i)}$, and denote the combined weights of OccamNet as $\mathbf{W}.$
@@
In principle, OccamLLM can work with any symbolic model, i.e., any model that can parameterize a set of symbolic functions or a distribution over such functions. We choose OccamNet as opposed to, for example, a transformer [31] or recurrent neural network [32], for two reasons: 1) OccamNet is interpretable, which we hypothesize makes controlling OccamNet an easier task for a decoder to learn, and 2) OccamNet is parallelizable over multiple samples, allowing for scalable training.
@@ 3.1.2 OccamLLM Decoder
The OccamLLM decoder takes the hidden states of a language model and outputs an initialization for OccamNet. This gives the LLM control over which function to apply on the inputs. The decoder acts on each input token separately, producing a different OccamNet initialization for each. Therefore, the arithmetic operations predicted may change along an input sequence, allowing OccamNet’s use for different computations in a single multi-token generation. This is crucial in multi-step reasoning scenarios where OccamNet is employed several times for different purposes.
Many decoder architectures are possible. We choose to parameterize the weights of each softmax layer of OccamNet independently, as $(\mathbf{W}^{(1)},\ldots,\mathbf{W}^{(l)})=(\text{Decoder}_{1}(\mathbf{h}),\ldots,\text{Decoder}_{l}(\mathbf{h})),$ where $\mathbf{h}$ are the hidden states of the language model. We choose
$\text{Decoder}_{i}(\mathbf{h})=\text{MLP}_{i}\left(\sum_{j=1}^{L}w_{i,j}\mathbf{h}_{j}\right)+\mathbf{W}^{*(i)}$ (1)
where $\mathbf{h}_{j}$ are the hidden states of the $j$th layer of the language model, $w_{i,j}$ are trainable weights, $\text{MLP}_{i}$ are two-layer multilayer perceptrons (MLPs), and $\mathbf{W}^{*(i)}$ are untrained weights which initialize all functions to have approximately equal probabilities according to the initialization scheme described in [11] and explained in Appendix D.4.
@@ 3.2 Data Generation
We create synthetic datasets to train the OccamLLM decoders, which contain instruction prompts for diverse arithmetic tasks. To generate datasets of arbitrary size, we create prompts with placeholders for numbers. Each prompt includes a question with number placeholders, the sampling value range for each number, and a function that computes the answer to the query given the sampled input numbers. The prompts fall into two main categories: purely arithmetic tasks and reasoning problems.
@@ 3.2.1 OccamNet Decoder Training Data
For training the decoder that controls the weights of OccamNet, we created two types of examples, single queries and concatenated queries. For single queries, we select a single prompt from the problems generated as discussed in Section 3.2. We use the Llama 3 8B Instruct chat template and fill in the query as the user input and the result as the assistant response, prepending “Answer = ” to the later in randomly selected samples (see Appendix A.1 for further details). For the concatenated queries of examples, we select a random number of prompts and concatenate the query-response pairs without using the Llama 3 8B Instruct chat template. The OccamNet decoder is trained to predict only the results of the last query in the sequence. This strategy helps OccamLLM to learn which operation to perform without becoming confused by earlier text, which is useful for continuous generation. To create the training dataset, each example is sampled by first randomly selecting whether to create a single or concatenated query, then randomly selecting the type(s) of prompt(s) used, and finally randomly sampling the input values from the range corresponding to each selected prompt.
@@ 3.2.2 OccamLLM Switch Training Data
To train the switch, we generate examples of possible LLM outputs for given input expressions, and label the outputs with sequences of 0s or 1s corresponding to whether the language model or the OccamNet output should be used for the next token. Some examples correspond to the prompts described in Section 3.2. For such examples, the LLM output is set to “The answer is” or “Answer = ” and the label sequence is all 0s with a 1 at the last token to indicate the system should use OccamNet only to compute the answer. We also manually created and labeled several other examples for diverse scenarios to explicitly teach the system in which cases it should or should not use OccamNet.
To create the training dataset, we concatenate a random number of the above user input - assistant output pairs in a conversational fashion, using the Llama 3 8B Instruct chat template.
@@ 3.3 OccamLLM Training
We train the OccamLLM decoder and the switch separately, as they do not share weights. In all cases, the weights of the LLM are kept frozen. In the first step, we train the system to predict the answer to examples generated by the method explained in Section 3.2.1. The OccamNet decoder processes the hidden states corresponding to the last token of the response and sets the weights of OccamNet such that the correct arithmetic expression is sampled. In this step, we use a rescaled REINFORCE [34] loss, which can also be interpreted as a Monte-Carlo estimate of the cross-entropy loss (see Appendix C.2):
@@
where $p_{W}[f]\equiv\text{ON}(f;\text{Decoder}_{W}(\mathbf{h}(x)))$ is the probability distribution represented by the decoder-initialized OccamNet.
Minimizing this loss steers the decoder towards assigning higher probabilities to the functions that maximize the reward $R(f(x),y)$, which measures the similarity between the correct answer $y$ and the prediction of OccamNet $f(x)$. We find setting $R(f(x),y)=1$ if $f(x)=y$, and 0 otherwise, most effective. We discuss the OccamNet loss in more detail in Appendix C.
@@ 4 Experiments
This single layer OccamNet can be invoked by the LLM several times during generation to perform complex arithmetic operations accurately. To use the trained OccamLlama for inference, we sample the highest probability function from OccamNet as described in Appendix D.3.